solve.
sqrt(y+4)=y-8

Answer by ankor@dixie-net.com(22740)

sqrt(y+4) = y - 8
:
Square both sides:
y + 4 = (y-8)^2
:
FOIL (y-8)(y-8)
y + 4 = y^2 - 16y + 64
:
0 = y^2 - 16y - y + 64 - 4
:
y^2 - 17y + 60 = 0; a quadratic equation:
:
Factors to:
(y-12)(y-5) = 0
:
y = 12
and
y = 5
:
Check both solutions in the original equation:
y = 12: sqrt(12+4) = 12 - 8; a good solution
:
y = 5: sqrt(5+4) = 5 - 8
+3 = - 3 not a good solution
:
:
Did this help?